40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 7 15:46:42 2013, Jenne, Update, Locking, PR2 flipped 
    Reply  Fri Feb 8 02:12:14 2013, Jenne, Update, Locking, PRMI work 
       Reply  Fri Feb 8 02:38:04 2013, Koji, Update, Locking, PRMI work 
       Reply  Fri Feb 8 11:01:18 2013, Jamie, Update, Locking, PRMI work 
          Reply  Fri Feb 8 12:39:32 2013, yuta, Update, Locking, PRMI work 
          Reply  Fri Feb 8 15:53:48 2013, Jenne, Update, Locking, PRMI work 
             Reply  Fri Feb 8 17:41:34 2013, Jenne, Update, Locking, PRMI work 
                Reply  Fri Feb 8 19:39:02 2013, Koji, Update, Locking, PRMI work 
                Reply  Fri Feb 8 20:05:15 2013, Jenne, Update, Locking, PRMI work 
Message ID: 8039     Entry time: Fri Feb 8 17:41:34 2013     In reply to: 8037     Reply to this: 8042   8043
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PRMI work 

 

 [Yuta, Jenne]

After much tweaking of the alignment using TT1, TT2 and PRM sliders, we were able to get a TEM00 mode locked with the half PRC!

PRCL gain is -0.010

FM4, 5 are always on.  FM2,3,6 (boosts and stack res-gains) are triggered to come on after the cavity is locked.

We see a little clipping of POPDC, even though there are 2 BSs in the beam path, to dump 50% and then 67% of the beam.  But it's not so much that we can't align. 

REFLDC goes from 28.5 to 24.5, so we don't have great visibility.

Please watch our awesome video of the cavity, where we demonstrate that the half cavity is stable:

The cavity is flashing for the 1st 15 sec, then locks.  Upper right is REFL, Lower right is POP, Upper left is back of the Faraday, Lower left is MC2F.   Note that we definitely see some not so beautiful modes flashing, but most of that is due to the half cavity length and thus greater degeneracy of modes.  Jamie is posting a HOM plot presently.

BEAM MOTION:

The beam is moving way more than it should be.  Right now the PRM oplev is not coming out of the vacuum, since the flat test mirror mount is obstructing it.  However, as we saw with other half-cavity tests, turning on the PRM oplev helps, but does not completely eliminate the beam motion.  We should consider putting oplevs on one of the passive TTs, at least temporarily, so we know what kind of motion is coming from where.

ELOG V3.1.3-